Can someone help me solve this problem? I am having a hard time with it, can you explain the steps?
Assume the profit earned by an artist in any given year is governed by the function P(n) = -2n^2 + 80n – 30 where n represents the number of paintings sold and P (in thousands of dollars) represents the profit. How many paintings should be sold to create a maximum profit? What is the maximum profit?
Thank you, Cynthia :)

Answer by richwmiller(17219)

You want the vertex.
The x component or n component will be the number of paints with maximum profit P(n)
what is the x component of the vertex -b/2a
-80/2*-2=80/4=20
plug 20 in for n and solve and that will give you the max of the parabola which should come out to 770
Everyone here wants help.
I generally avoid the problems with begging.